OpenStudy (anonymous):

24, 53, 38, 12, 31, 19, 26 find the range, median, first and third quartiles, and interquartile range

OpenStudy (anonymous):

@jabez177

OpenStudy (anonymous):

@jigglypuff314

OpenStudy (anonymous):

@phi

Directrix (directrix):

I will help you with this is you post the data points in ascending order (low to high). Thanks. @JessicaH123

OpenStudy (anonymous):

12, 19, 24, 26, 31, 38, 53

Directrix (directrix):

12, 19, 24, 26, 31, 38, 53 The range of the data set is the difference between the largest value of the data set and the smallest value of the data set. What do you think that will be?

OpenStudy (anonymous):

53-12 = 41

OpenStudy (anonymous):

?

Directrix (directrix):

Correct.

alones (alones):

You got it right! way to go <3

Directrix (directrix):

median - the score in the middle of the set of scores when the scores are in order 12, 19, 24, 26, 31, 38, 53 Do you see it above? The median is also known as the second quartile.

OpenStudy (anonymous):

26?

Directrix (directrix):

Yes

OpenStudy (anonymous):

thanks @AloneS

Directrix (directrix):

The lower quartile Q1 is the median of just the scores to the left of Q2, the median of the entire data set. 12, 19, 24, Do you see Q1 ?

OpenStudy (anonymous):

19

Directrix (directrix):

Correct. The upper quartile Q3 is the median of the data points to the right of the median of the entire data set. 31, 38, 53 Do you see the upper quartile?

OpenStudy (anonymous):

38

Directrix (directrix):

Correct. interquartile range The IQR (Interquartile Range) is the difference between Q3 and Q1. We already have those values so all we have to do is to subtract. Q3 - Q1 = ?

Directrix (directrix):

Try again. Q3 = 38 Q1 = 19 Q3 - Q1 = ?

OpenStudy (anonymous):

oh sorry i was looking at another problem

OpenStudy (anonymous):

19

Directrix (directrix):

Correct.

OpenStudy (anonymous):

thanks

Directrix (directrix):

You are welcome.

OpenStudy (anonymous):

could you have another problem this one is so confused?

Directrix (directrix):

Close this thread and post it in a new one I will try to help, okay?

OpenStudy (anonymous):

oh

Directrix (directrix):

Is something wrong?

OpenStudy (anonymous):

no